if ABCD is similar to EFGD then x=
                  \  A___________5___________B
                 /   /                       \
                /   /   Y                     \
               9  E/_______________F           \ 7.5
              /   /                \            \
             /   /                  \x           \
            /   /                    \            \ 
           \ D_/______________________\G___________\C

the 9 is on the outside of AED

Answer by Cromlix(4381)

Hi there,
To answer for the value of 'x'
you need to include the size of
one of the sides of EFGD.
Then you can you can set up an equation:
Say you have the size of EF.
AB/EF = BC/FG
5/EF = 7.5/x
Then you can cross multiply
5*x = 7.5*EF
x = (7.5 * EF)/5
You put in the value of EF and multiply it by 7.5
Then divide by 5 to obtain the value of x
Hope this helps :-)